# Delving Deep into Diffusion Transformers for Image and Video Generation

### v1

@@ Abstract
In this study, we explore Transformer-based diffusion models for image and video generation. Despite the dominance of Transformer architectures in various fields due to their flexibility and scalability, the visual generative domain primarily utilizes CNN-based U-Net architectures, particularly in diffusion-based models. We introduce GenTron, a family of Generative models employing Transformer-based diffusion, to address this gap. Our initial step was to adapt Diffusion Transformers (DiTs) from class to text conditioning, a process involving thorough empirical exploration of the conditioning mechanism. We then scale GenTron from approximately 900M to over 3B parameters, observing significant improvements in visual quality. Furthermore, we extend GenTron to text-to-video generation, incorporating novel motion-free guidance to enhance video quality. 1https://www.shoufachen.com/gentron_website/ In human evaluations against SDXL, GenTron achieves a 51.1% win rate in visual quality (with a 19.8% draw rate), and a 42.3% win rate in text alignment (with a 42.9% draw rate). GenTron also excels in the T2I-CompBench, underscoring its strengths in compositional generation. We believe this work will provide meaningful insights and serve as a valuable reference for future research. The website of GenTron is available1.
@@ 1 Introduction
Diffusion models have recently shown remarkable progress in content creation, impacting areas such as image generation [25, 53, 51], video production [26, 5, 57], audio synthesis [54], and code generation [58]. Among these notable developments, the convolutional U-Net architecture has emerged as the predominant backbone design, a choice that stands in contrast to the prevailing trend in natural language processing [17, 7, 60] and computer visual perception [18, 34, 67] domains, where attention-based transformer architectures [61] have empowered a renaissance and become increasingly dominant. To provide a comprehensive understanding of Transformers in diffusion generation and to bridge the gap in architectural choices between visual generation and the other two domains — visual perception and NLP — a thorough investigation of visual generation using Transformers is of substantial scientific value.
We focus on diffusion models with Transformers in this work. Specifically, our starting point is the foundational work known as DiT [41], which introduced a class-conditioned latent diffusion model that employs a Transformer to replace the traditionally used U-Net architecture. We first overcome the limitation of the original DiT model, which is constrained to handling only a restricted number (e.g., 1000) of predefined classes, by utilizing language embeddings derived from open-world, free-form text captions instead of predefined one-hot class embeddings. Along the way, we comprehensively investigate conditioning strategies, including (1) conditioning architectures: adaptive layer norm (adaLN) [42] vs. cross-attention [61]; and (2) text encoding methods: a generic large language model [12] vs. the language tower of multimodal models [47], or the combination of both of them. We additionally carry out comparative experiments and offer detailed empirical analyses to evaluate the effectiveness of these conditioning strategies.
@@ 2 Related Work
Recently, Transformer-based Diffusion models have attracted increasing research interest. Among these, U-ViT [3] treats all inputs as tokens by integrating transformer blocks with a U-net architecture. In contrast, DiT [41] employs a simpler, non-hierarchical transformer structure. MDT [20] and MaskDiT [69] enhance DiT’s training efficiency by incorporating the mask strategy [22]. Dolfin [64] is a transformer-based model for layout generation. Concurrently to this work, PixArt-$\alpha$ [9] demonstrates promising outcomes in Transformer-based T2I diffusion. It’s trained using a three-stage decomposition process with high-quality data. Our work diverges from PixArt-$\alpha$ in key aspects. Firstly, while PixArt-$\alpha$ emphasizes training efficiency, our focus is on the design choice of conditioning strategy and scalability in T2I Transformer diffusion models. Secondly, we extend our exploration beyond image generation to video diffusion. We propose an innovative approach in video domain, which is covered by PixArt-$\alpha$.
@@ 3.3.2 Motion-Free Guidance
Model Attribute Binding Object Relationship Complex Mean
Color Shape Texture Spatial Non-spatial
@@ 4.1 Implementation Details
[…]
@@ 4.2 Main Results of GenTron-T2I
Figure 6: GenTron-T2V results. Prompts are (left-to-right, top-to-bottom): “Teddy bear walking down 5th Avenue front view beautiful sunset”, “A dog swimming”, “A giant tortoise is making its way across the beach”, and “A dolphin jumping out of the water”.
@@
In Table 3, we showcase the alignment evaluation results from T2I-CompBench. Our method demonstrates outstanding performance in all areas, including attribute binding, object relationships, and complex compositions. This indicates a heightened proficiency in compositional generation, with a notable strength in color binding. In this aspect, our approach surpasses the previous state-of-the-art (SoTA) benchmark set  [9] by over 7%.
@@
We visualize the human preference of our method versus Latent Diffusion XL in Fig. 7. We used standard prompts in PartiPrompt2 [66] to generate 100 images using both methods and ask people for their preference blindly after shuffling. We received a total of three thousand responses on the comparisons of visual quality and text faithfulness, with our method emerging as the clear winner by a significant margin.

### v2

# GenTron: Diffusion Transformers for Image and Video Generation

@@ Abstract
In this study, we explore Transformer-based diffusion models for image and video generation. Despite the dominance of Transformer architectures in various fields due to their flexibility and scalability, the visual generative domain primarily utilizes CNN-based U-Net architectures, particularly in diffusion-based models. We introduce GenTron, a family of Generative models employing Transformer-based diffusion, to address this gap. Our initial step was to adapt Diffusion Transformers (DiTs) from class to text conditioning, a process involving thorough empirical exploration of the conditioning mechanism. We then scale GenTron from approximately 900M to over 3B parameters, observing improvements in visual quality. Furthermore, we extend GenTron to text-to-video generation, incorporating novel motion-free guidance to enhance video quality. 1https://www.shoufachen.com/gentron_website/ In human evaluations against SDXL, GenTron achieves a 51.1% win rate in visual quality (with a 19.8% draw rate), and a 42.3% win rate in text alignment (with a 42.9% draw rate). GenTron notably performs well in T2I-CompBench, highlighting its compositional generation ability. We hope GenTron could provide meaningful insights and serve as a valuable reference for future research. Please refer to the website1 and the arXiv version for the most up-to-date results: https://arxiv.org/abs/2312.04557.
@@ 1 Introduction
Diffusion models have recently shown remarkable progress in content creation, impacting areas such as image generation [27, 57, 55], video generation [28, 5, 60, 7], and editing [34, 14]. Among these notable developments, the CNN-based U-Net architecture has emerged as the predominant backbone design, a choice that stands in contrast to the prevailing trend in natural language processing [19, 8, 62] and computer visual perception [20, 38, 69] domains, where attention-based transformer architectures [63] have empowered a renaissance and become increasingly dominant. To provide a comprehensive understanding of Transformers in diffusion generation and to bridge the gap in architectural choices between visual generation and the other two domains — visual perception and NLP — a thorough investigation of visual generation using Transformers is of substantial scientific value.
We focus on diffusion models with Transformers in this work. Specifically, our starting point is the foundational work known as DiT [45], which introduced a class-conditioned latent diffusion model that employs a Transformer to replace the traditionally used U-Net architecture. We overcome the limitation of the original DiT model, which is constrained to handling only a restricted number (e.g., 1000) of predefined classes, by utilizing language embeddings derived from open-world, free-form text captions instead of predefined one-hot class embeddings. Along the way, we comprehensively investigate conditioning strategies, including (1) conditioning architectures: adaptive layer norm (adaLN) [46] vs. cross-attention [63]; and (2) text encoding methods: a generic large language model [13] vs. the language tower of multimodal models [51], or the combination of both of them. We additionally carry out comparative experiments and offer detailed empirical analyses to evaluate the effectiveness of these conditioning strategies.
@@ 2 Related Work
Recently, Transformer-based Diffusion models have attracted increasing research interest. Among these, U-ViT [3] treats all inputs as tokens by integrating transformer blocks with a U-net architecture. In contrast, DiT [45] employs a simpler, non-hierarchical transformer structure. MDT [22] and MaskDiT [71] enhance DiT’s training efficiency by incorporating the mask strategy [24]. Dolfin [66] is a transformer-based model for layout generation. Concurrently to this work, PixArt-$\alpha$ [10] demonstrates promising outcomes in Transformer-based T2I diffusion. It’s trained using a three-stage decomposition process with high-quality data. Our work diverges from PixArt-$\alpha$ in key aspects. Firstly, while PixArt-$\alpha$ emphasizes training efficiency, our focus is on the design choice of conditioning strategy and scalability in T2I Transformer diffusion models. Secondly, we extend our exploration beyond image generation to video diffusion. We propose an innovative approach in video domain, which is not covered by PixArt-$\alpha$.
@@ 3.3.2 Motion-Free Guidance
Model Attribute Binding Obj. Relation Comp. Mean
Color Shape Texture Spat. Non-spat.
@@ 4.1 Implementation Details
For all GenTron model variations, we employ the AdamW [39] optimizer, maintaining a constant learning rate of 1$\times$10-4. We train our T2I GenTron models in a multi-stage procedure [55, 47] with an internal dataset, including a low-resolution (256$\times$256) training with a batch size of 2048 and 500K optimization steps, as well as high-resolution (512$\times$512) with a batch size of 784 and 300K steps. For the GenTron-G/2 model, we further integrate Fully Sharded Data Parallel (FSDP) [70] and activation checkpointing (AC), strategies specifically adopted to optimize GPU memory usage. In our video experiments, we train videos on a video dataset that comprises approximately 34M videos. To optimize storage usage and enhance data loading efficiency, the videos are pre-processed to a resolution with a short side of 512 pixels and a frame rate of 24 FPS. We process batches of 128 video clips. Each clip comprises 8 frames, captured at a sampling rate of 4 FPS.
@@ 4.2 Main Results of GenTron-T2I
Figure 6: GenTron-T2V examples. Prompts are “A giant tortoise is making its way across the beach” and “A dog swimming”.
@@
[…]
Imagen 3.0B 7.27 0.27 -
Parti-750M 0.8B 10.71 - -
Parti-3B 3.0B 8.10 - -
GigaGAN 1.0B 9.09 0.322 -
MUSE-3B 3.0B 7.88 0.320 -
SD v1.4 0.9B 12.94 0.325 31.50
SDXL 2.6B 17.82 0.329 44.41
GenTron-XL/2 0.9B 14.21 0.326 49.13
GenTron-G/2 3.1B 14.53 0.335 49.99
Table 4: Comparison to T2I models.
@@
In Table 3, we showcase the alignment evaluation results from T2I-CompBench. Our method demonstrates outstanding performance in all areas, including attribute binding, object relationships, and complex compositions. This indicates a heightened proficiency in compositional generation, with a notable strength in color binding. In this aspect, our approach surpasses the previous state-of-the-art (SoTA) work [10] by over 7%.
In Table 4, we further thoroughly compare the zero-shot FID-30K, CLIP-Score, and T2I-CompBench metrics with previous T2I models. In contrast to U-Net based SDv1.4, GenTron uses about four times less data (550M vs. 2B), while achieving better CLIP-scores and T2I-CompBench results. Although GenTron does not achieve a superior FID score compared to SDv1.4, it’s important to note that recent studies, including Pick-a-Pic [37] and SDXL [47], have highlighted that FID scores often misrepresent human preferences in generative models, sometimes even negatively correlating with visual aesthetics.
@@
Figure 7 shows the human preference of our method versus SDXL. We used standard prompts in PartiPrompt2 [68] to generate 100 images using both methods and ask people for their preference blindly after shuffling. We received a total of three thousand responses on the comparisons of visual quality and text faithfulness, with our method emerging as the clear winner by a clear margin.
@@ 5 Conclusion
Acknowledgement. This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000 and the General Research Fund of Hong Kong No.17200622.
@@ Appendix C Additional GenTron-T2I Examples
Appendix D Additional GenTron-T2V Examples
Additional GenTron-T2V results are available on our website1.